OpenStudy (anonymous):

Jaime knows that the area of a triangle can be found by the formula A = 1/2bh. If the area of a triangle is 26 m2 and the height is 13 meters, what is the length of the base? 2 meters 8 meters 4 meters 1 meter

OpenStudy (anonymous):

please help i need a awnserr.

OpenStudy (anonymous):

@Abmon98

OpenStudy (abmon98):

26=1/2*b*13 26=13b/2 13b=26*2 b=26*2/13=4meter

OpenStudy (anonymous):

thanks so its 4

OpenStudy (abmon98):

yes

OpenStudy (anonymous):

Solve for q if 8/q = 64/16 4 2 8 16

OpenStudy (anonymous):

i dont get this one.

OpenStudy (anonymous):

@Abmon98

OpenStudy (abmon98):

post a new question dont post your question in your old one